Determine the value of the equilibrium constant, Q for the chemical reaction given below, if 0.120 moles of Nitrogen (Ni), 0.077 moles of sulfate (SO4), hydrogen (H) 0.0104 and 0.0084 moles of hot water are there in a 4 liters cylinder and at an equilibrium.


Ni= 0.120


SO4= 0.077


H= 0.0104


hot water (H2O)= 0.0084


cylinder= 4 liters




substitute the value of the compounds


1
Expert's answer
2021-04-05T07:49:20-0400

Δ[NH3]=−2×Δ[N2]=−2×(0.11M)=−0.22MΔ